# T2V-Turbo: Breaking the Quality Bottleneck of Video Consistency Model with Mixed Reward Feedback

### v1

@@ Abstract
Figure 1: By integrating reward feedback during consistency distillation from VideoCrafter2 (Chen et al., 2024), our T2V-Turbo (VC2) can generate high-quality videos with 4-8 inference steps, breaking the quality bottleneck of a VCM (Wang et al., 2023a).
@@ 1 Introduction
In this paper, we aim to achieve fast and high-quality video generation by breaking the quality bottleneck of a VCM. We introduce T2V-Turbo, which integrates reward feedback from a mixture of RMs into the process of distilling a VCM from a teacher T2V model. Besides utilizing an image-text RM to align individual video frames with human preference, we further incorporate reward feedback from a video-text RM to comprehensively evaluate the temporal dynamics and transitions in the generated videos. We highlight that our reward optimization avoids tackling the highly memory-intensive issues associated with backpropagating gradients through an iterative sampling process. Instead, we directly optimize rewards of the single-step generations that arise from computing the CD loss, effectively bypassing the memory constraints faced by conventional methods that optimize a DM (Yuan et al., 2023; Xu et al., 2024; Clark et al., 2023; Prabhudesai et al., 2023).
@@ 2 Preliminaries
where $c_{\text{skip}}(t)$ and $c_{\text{out}}(t)$ are differentiable functions with $c_{\text{skip}}(\epsilon)=1$ and $c_{\text{out}}(\epsilon)=0$, and $F_{\theta}$ is modeled as a neural network. We can distill a $\boldsymbol{f}_{\theta}$ from a pre-trained T2V DM by minimizing the consistency distillation (CD) Song et al. (2023); Luo et al. (2023a) loss as below
@@
We follow the LCM paper (Luo et al., 2023a) to use DDIM Song et al. (2020b) as the ODE solver $\Psi$ and defer the formula of the DDIM solver to Appendix A.
@@ 3 Training T2V-Turbo with Mixed Reward Feedback
In this section, we present the training pipeline to derive our T2V-Turbo. To facilitate fast and high-quality video generation, we integrate reward feedback from multiple RMs into the LCD process when distilling from a teacher T2V model. Figure 2 provides an overview of our framework. Notably, we directly leverage the single-step generation $\hat{\mathbf{z}}_{0}=\boldsymbol{f}_{\theta}\left(\mathbf{z}_{t_{n+k}},\omega,\mathbf{c},t_{n+k}\right)$ arise from computing the CD loss $L_{\text{CD}}$ (6) and optimize the video $\hat{\mathbf{x}}_{0}=\mathcal{D}(\hat{\mathbf{z}}_{0})$ decoded from it towards multiple differentiable RMs. As a result, we avoid the challenges associated with backpropagating gradients through an iterative sampling process, which is often confronted by conventional methods optimizing DMs (Clark et al., 2023; Xu et al., 2024; Yuan et al., 2023).
@@ 3.2 Optimizing Video-Text Feedback Model
Existing image-text RMs (Wu et al., 2023a; Xu et al., 2024; Kirstain et al., 2024) are limited to assessing the alignment between individual video frames and the text prompt and thus cannot evaluate through the temporal dimensions that involve inter-frame dependencies, such as motion dynamic and transitions (Huang et al., 2024; Liu et al., 2023). To address these shortcomings, we further leverage a video-text RM $\mathcal{R}_{\text{vid}}$ to assess the generated videos. The corresponding objective $J_{\text{vid}}$ is given below
@@ 4.2 Human Evaluation with 700 EvalCrafter Prompts
To verify the effectiveness of our T2V-Turbo, we compare the 4-step and 8-step generations from our T2V-Turbo with the 50-step DDIM samples from the corresponding teacher T2V models. We further compare the 4-step generations between our T2V-Turbo and their baseline VCMs when distilled from the same teacher T2V model.
We leverage the 700 prompts from the EvalCrafter (Liu et al., 2023) video evaluation benchmark, which are constructed based on real-world user data. We hire human annotators from Amazon Mechanical Turk to compare videos generated from different models given the same prompt. For each comparison, the annotators need to answer three questions: Q1) Which video is more visually appealing? Q2) Which video better fits the text description? Q3) Which video do you prefer given the prompt? Appendix D includes additional details about how we setup the human evaluations.
Figure 3 provides the full human evaluation results. We also qualitatively compare different methods in Figure 4. Appendix F further includes additional qualitative comparison results. Notably, the 4-step generations from our T2V-Turbo are favored by humans over the 50-step generation from their teacher T2V model, representing a 25 times inference acceleration with improving performance. By increasing the inference steps to 8, we can further improve the visual quality and text-video alignment of videos generated from our T2V-Turbo, reflected by the fact that our 8-step generations are more likely to be favored by the human compared to our 4-step generations in terms of all 3 evaluated metrics. Additionally, our T2V-Turbo significantly outperforms its baseline VCM, demonstrating the effectiveness of methods that incorporate a mixture of reward feedback into the model training.
@@ 4.3 Ablation Studies
Ablating RMs $\mathcal{R}_{\text{img}}$ and $\mathcal{R}_{\text{vid}}$. Recall that the training of our T2V-Turbo incorporate reward feedback from both $\mathcal{R}_{\text{img}}$ and $\mathcal{R}_{\text{vid}}$. To demonstrate the effectiveness of each individual RM, we perform ablation study by training VCM (VC2) + $\mathcal{R}_{\text{vid}}$ and VCM (VC2) + $\mathcal{R}_{\text{img}}$, which only incorporate feedback from $\mathcal{R}_{\text{vid}}$ and $\mathcal{R}_{\text{img}}$, respectively. Again, we evaluate the 4-step generations from different methods on VBench. Results in Table 2 show that incorporating feedback from either $\mathcal{R}_{\text{img}}$ or $\mathcal{R}_{\text{vid}}$ leads to performance improvement over the baseline VCM. Notably, optimizing $\mathcal{R}_{\text{img}}$ alone can already lead to substantial performance gains, while incorporating feedback from $\mathcal{R}_{\text{vid}}$ can further improve the Semantic Score on VBench, leading to better text-video alignment.
@@ 5 Related Work
Accelerating inference of Diffusion Models. Various methods have been proposed to accelerate the sampling process of a DM, including advanced numerical ODE solvers Song et al. (2020b); Lu et al. (2022a, b); Zheng et al. (2022); Dockhorn et al. (2022); Jolicoeur-Martineau et al. (2021) and distillation techniques Luhman and Luhman (2021); Salimans and Ho (2021); Meng et al. (2023); Zheng et al. (2023). Recently, Consistency Model (Song et al., 2023; Luo et al., 2023a) is proposed to facilitate fast inference by learning a consistency function to map any point at the ODE trajectory to the origin. Li et al. (2024) proposes to augment consistency distillation with an objective to optimize image-text RM to achieve fast and high-quality image generation. Our work extends it for T2V generation, incorporating reward feedback from both an image-text RM and a video-text RM.
Vision-and-language Reward Models. There have been various open-sourced image-text RMs that are trained to mirror human preferences given a text-image pair, including HPS (Wu et al., 2023b, a), ImageReward Xu et al. (2024), and PickScore Kirstain et al. (2024), which are obtained by finetuning a image-text foundation model such as CLIP Radford et al. (2021) and BLIP Li et al. (2022), on human preference data. However, to the best of our knowledge, no video-text RMs, e.g., T2VScore (Wu et al., 2024), that mirrors human preference on a text-video pair has been released to the public. In this paper, we choose HPSv2.1 as our image-text RM and directly employ the video foundation models ViCLIP (Wang et al., 2023d) and InterVid S2 (Wang et al., 2024) that are trained for general video-text understanding as our video-text RM. Empirically, we show that incorporating feedback from these RMs can improve the performance of our T2V-Turbo.
Learning from Human/AI Feedback has been proven as an effective way to align the output from a generative model with human preference (Leike et al., 2018; Ziegler et al., 2019; Ouyang et al., 2022; Stiennon et al., 2020; Rafailov et al., 2024). In the field of image generation, various methods have been proposed to align a text-to-image model with human preference, including RL based methods Fan et al. (2024); Prabhudesai et al. (2023); Zhang et al. (2024) and backpropagation-based reward finetuning methods Clark et al. (2023); Xu et al. (2024); Prabhudesai et al. (2023). Recently, InstructVideo (Yuan et al., 2023) extends the reward-finetuning methods to optimize a T2V model. However, it still employs an image-text RM to provide reward feedback without considering the transition dynamic of the generated video. In contrast, our work incorporates reward feedback from both an image-text and video-text RM, providing comprehensive feedback to our T2V-Turbo.
@@ Appendix A Experiment and Hyperparameter (HP) Details
As mentioned in Sec. 2, we employ the DDIM Song et al. [2020b] ODE solver $\Psi_{\text{DDIM}}$ by following the practice of Luo et al. [2023a]. Its formula from $t_{n+k}$ to $t_{n}$ is given below
@@
where $\hat{\boldsymbol{\epsilon}}_{\psi}$ denotes the noise prediction model from the teacher T2V model. We refer interested readers to the original LCM paper Luo et al. [2023a] for further details.
@@ Appendix B Psudoe-codes for Training our T2V-Turbo
We include the pseudo-codes for training our T2V-Turbo in Algorithm 1. We use the red color to highlight the difference from the standard (latent) consistency distillation Luo et al. [2023a], Song et al. [2023].
@@ Appendix C Further Details about VBench
  • Motion Smoothness (Motion Smooth.) is evaluated by motion priors in the video frame interpolation model [Li et al., 2023a].
@@
  • Human Action is evaluated by the UMT model [Li et al., 2023b].
@@ Appendix D Human Evaluation Details
As mentioned in Sec. 4.2, we hire labelers from Amazon Mechanical Turk platform and form the video comparison task as many batches of HITs. Specifically, we choose lablers from English-speaking countries, including AU, CA, NZ, GB, and US. Each task needs around 30 seconds to complete, and we pay each submitted HIT with 0.2 US dollars. Therefore, the hourly payment is about 24 US dollars.
@@ Appendix E Additional Ablation Studies
In this section, we provide the full ablation results performed in Table 3, which can be found in Table 5. We further examine the effect of different choices of $\mathcal{R}_{\text{img}}$. In the initial stage of our project, we train VCM (VC2) + $\mathcal{R}_{\text{img}}$ with several different image-text RMs, including HPSv2.1 [Wu et al., 2023a], PickScore [Kirstain et al., 2024], and ImageReward [Xu et al., 2024]. We collect the 4-step generations from each method and qualitatively compare them with the 4-step generation from the baseline VCM. As shown in Figure 6, incorporating reward feedback from any of these $\mathcal{R}_{\text{img}}$ leads to quality improvment over the baseline VCM (VC2).

### v2

@@ Abstract
Figure 1: By integrating reward feedback during consistency distillation from VideoCrafter2 (Chen et al., 2024), our T2V-Turbo (VC2) can generate high-quality videos with 4-8 inference steps, breaking the quality bottleneck of a VCM (Wang et al., 2023a). Appendix F includes the corresponding text prompts.
@@ 1 Introduction
In this paper, we aim to achieve fast and high-quality video generation by breaking the quality bottleneck of a VCM. We introduce T2V-Turbo, which integrates reward feedback from a mixture of RMs into the process of distilling a VCM from a teacher T2V model. Besides utilizing an image-text RM to align individual video frames with human preference, we further incorporate reward feedback from a video-text RM to comprehensively evaluate the temporal dynamics and transitions in the generated videos. We highlight that our reward optimization avoids tackling the highly memory-intensive issues associated with backpropagating gradients through an iterative sampling process. Instead, we directly optimize rewards of the single-step generations that arise from computing the CD loss, effectively bypassing the memory constraints faced by conventional methods that optimize a DM (Yuan et al., 2023; Xu et al., 2024a; Clark et al., 2023; Prabhudesai et al., 2023).
@@ 2 Preliminaries
where $c_{\text{skip}}(t)$ and $c_{\text{out}}(t)$ are differentiable functions with $c_{\text{skip}}(\epsilon)=1$ and $c_{\text{out}}(\epsilon)=0$, and $F_{\theta}$ is modeled as a neural network. We can distill a $\boldsymbol{f}_{\theta}$ from a pre-trained T2V DM by minimizing the consistency distillation (CD) (Song et al., 2023; Luo et al., 2023a) loss as below
@@
We follow the LCM paper (Luo et al., 2023a) to use DDIM (Song et al., 2020b) as the ODE solver $\Psi$ and defer the formula of the DDIM solver to Appendix A.
@@ 3 Training T2V-Turbo with Mixed Reward Feedback
In this section, we present the training pipeline to derive our T2V-Turbo. To facilitate fast and high-quality video generation, we integrate reward feedback from multiple RMs into the LCD process when distilling from a teacher T2V model. Figure 2 provides an overview of our framework. Notably, we directly leverage the single-step generation $\hat{\mathbf{z}}_{0}=\boldsymbol{f}_{\theta}\left(\mathbf{z}_{t_{n+k}},\omega,\mathbf{c},t_{n+k}\right)$ arise from computing the CD loss $L_{\text{CD}}$ (6) and optimize the video $\hat{\mathbf{x}}_{0}=\mathcal{D}(\hat{\mathbf{z}}_{0})$ decoded from it towards multiple differentiable RMs. As a result, we avoid the challenges associated with backpropagating gradients through an iterative sampling process, which is often confronted by conventional methods optimizing DMs (Clark et al., 2023; Xu et al., 2024a; Yuan et al., 2023).
@@ 3.2 Optimizing Video-Text Feedback Model
Existing image-text RMs (Wu et al., 2023a; Xu et al., 2024a; Kirstain et al., 2024) are limited to assessing the alignment between individual video frames and the text prompt and thus cannot evaluate through the temporal dimensions that involve inter-frame dependencies, such as motion dynamic and transitions (Huang et al., 2024; Liu et al., 2023). To address these shortcomings, we further leverage a video-text RM $\mathcal{R}_{\text{vid}}$ to assess the generated videos. The corresponding objective $J_{\text{vid}}$ is given below
@@ 4.2 Human Evaluation with 700 EvalCrafter Prompts
To verify the effectiveness of our T2V-Turbo, we compare the 4-step and 8-step generations from our T2V-Turbo with the 50-step DDIM samples from the corresponding teacher T2V models. We further compare the 4-step generations between our T2V-Turbo and their baseline VCMs when distilled from the same teacher T2V model. We leverage the 700 prompts from the EvalCrafter (Liu et al., 2023) video evaluation benchmark, which are constructed based on real-world user data.
We hire human annotators from Amazon Mechanical Turk to compare videos generated from different models given the same prompt. For each comparison, the annotators need to answer three questions: Q1) Which video is more visually appealing? Q2) Which video better fits the text description? Q3) Which video do you prefer given the prompt? Appendix D includes additional details about how we set up the human evaluations.
Figure 3 provides the full human evaluation results. We also qualitatively compare different methods in Figure 4. Due to limited space, we include additional qualitative comparison results in Appendix F. Notably, the 4-step generations from our T2V-Turbo are favored by humans over the 50-step generation from their teacher T2V model, representing a 12.5 times inference acceleration with improving performance. By increasing the inference steps to 8, we can further improve the visual quality and text-video alignment of videos generated from our T2V-Turbo, reflected by the fact that our 8-step generations are more likely to be favored by the human compared to our 4-step generations in terms of all 3 evaluated metrics. Additionally, our T2V-Turbo significantly outperforms its baseline VCM, demonstrating the effectiveness of our methods, which incorporate a mixture of reward feedback into the model training.
@@ 4.3 Ablation Studies
Ablating RMs $\mathcal{R}_{\text{img}}$ and $\mathcal{R}_{\text{vid}}$. Recall that the training of our T2V-Turbo incorporate reward feedback from both $\mathcal{R}_{\text{img}}$ and $\mathcal{R}_{\text{vid}}$. To demonstrate the effectiveness of each individual RM, we perform ablation study by training VCM (VC2) + $\mathcal{R}_{\text{vid}}$ and VCM (VC2) + $\mathcal{R}_{\text{img}}$, which only incorporate feedback from $\mathcal{R}_{\text{vid}}$ and $\mathcal{R}_{\text{img}}$, respectively. Again, we evaluate the 4-step generations from different methods on VBench. Results in Table 2 show that incorporating feedback from either $\mathcal{R}_{\text{img}}$ or $\mathcal{R}_{\text{vid}}$ leads to performance improvement over the baseline VCM. Notably, optimizing $\mathcal{R}_{\text{img}}$ alone can already lead to substantial performance gains, while incorporating feedback from $\mathcal{R}_{\text{vid}}$ can further improve the Semantic Score on VBench, leading to better text-video alignment. In Appendix H, we qualitatively compare the videos generated by our T2V-Turbo and VCM + $\mathcal{R}_{\text{img}}$, corroborating the effectiveness of our mixture of RMs design.
@@ 5 Related Work
Accelerating inference of Diffusion Models. Various methods have been proposed to accelerate the sampling process of a DM, including advanced numerical ODE solvers (Song et al., 2020b; Lu et al., 2022a, b; Zheng et al., 2022; Dockhorn et al., 2022; Jolicoeur-Martineau et al., 2021) and distillation techniques (Luhman and Luhman, 2021; Salimans and Ho, 2021; Meng et al., 2023; Zheng et al., 2023). Recently, Consistency Model (Song et al., 2023; Luo et al., 2023a) is proposed to facilitate fast inference by learning a consistency function to map any point at the ODE trajectory to the origin. Li et al. (2024) proposes to augment consistency distillation with an objective to optimize image-text RM to achieve fast and high-quality image generation. Our work extends it for T2V generation, incorporating reward feedback from both an image-text RM and a video-text RM.
Vision-and-language Reward Models. There have been various open-sourced image-text RMs that are trained to mirror human preferences given a text-image pair, including HPS (Wu et al., 2023b, a), ImageReward (Xu et al., 2024a), and PickScore (Kirstain et al., 2024), which are obtained by finetuning a image-text foundation model such as CLIP (Radford et al., 2021) and BLIP (Li et al., 2022), on human preference data. However, to the best of our knowledge, no video-text RMs, e.g., T2VScore (Wu et al., 2024), that mirrors human preference on a text-video pair has been released to the public. In this paper, we choose HPSv2.1 as our image-text RM and directly employ the video foundation models ViCLIP (Wang et al., 2023d) and InterVid S2 (Wang et al., 2024) that are trained for general video-text understanding as our video-text RM. Empirically, we show that incorporating feedback from these RMs can improve the performance of our T2V-Turbo.
Learning from Human/AI Feedback has been proven as an effective way to align the output from a generative model with human preference (Leike et al., 2018; Ziegler et al., 2019; Ouyang et al., 2022; Stiennon et al., 2020; Rafailov et al., 2024; Xu et al., 2024b). In the field of image generation, various methods have been proposed to align a text-to-image model with human preference, including RL (Sutton and Barto, 2018; Li et al., 2020, 2023a) based methods (Fan et al., 2024; Prabhudesai et al., 2023; Zhang et al., 2024) and backpropagation-based reward finetuning methods (Clark et al., 2023; Xu et al., 2024a; Prabhudesai et al., 2023). Recently, InstructVideo (Yuan et al., 2023) extends the reward-finetuning methods to optimize a T2V model. However, it still employs an image-text RM to provide reward feedback without considering the transition dynamic of the generated video. In contrast, our work incorporates reward feedback from both an image-text and video-text RM, providing comprehensive feedback to our T2V-Turbo.
@@ 6 Conclusion and Limitations
Acknowledgement
The work was funded by an unrestricted gift from Google, and we are grateful for their generous sponsorship. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the sponsors’ official policy, expressed or inferred.
@@ Appendix A Experiment and Hyperparameter (HP) Details
As mentioned in Sec. 2, we employ the DDIM [Song et al., 2020b] ODE solver $\Psi_{\text{DDIM}}$ by following the practice of Luo et al. [2023a]. Its formula from $t_{n+k}$ to $t_{n}$ is given below
@@
where $\hat{\boldsymbol{\epsilon}}_{\psi}$ denotes the noise prediction model from the teacher T2V model. We refer interested readers to the original LCM paper [Luo et al., 2023a] for further details.
@@ Appendix B Psudoe-codes for Training our T2V-Turbo
We include the pseudo-codes for training our T2V-Turbo in Algorithm 1. We use the red color to highlight the difference from the standard (latent) consistency distillation [Luo et al., 2023a, Song et al., 2023].
@@ Appendix C Further Details about VBench
  • Motion Smoothness (Motion Smooth.) is evaluated by motion priors in the video frame interpolation model [Li et al., 2023b].
@@
  • Human Action is evaluated by the UMT model [Li et al., 2023c].
@@ Appendix D Human Evaluation Details
As mentioned in Sec. 4.2, we hire labelers from Amazon Mechanical Turk platform and form the video comparison task as many batches of HITs. Specifically, we choose labelers from English-speaking countries, including AU, CA, NZ, GB, and the US. Each task needs around 30 seconds to complete, and we pay each submitted HIT with 0.2 US dollars. Therefore, the hourly payment is about 24 US dollars.
@@ Appendix E Additional Ablation Studies
In this section, we provide the full ablation results performed in Table 3, which can be found in Table 5. We further examine the effect of different choices of $\mathcal{R}_{\text{img}}$. In the initial stage of our project, we train VCM (VC2) + $\mathcal{R}_{\text{img}}$ with several different image-text RMs, including HPSv2.1 [Wu et al., 2023a], PickScore [Kirstain et al., 2024], and ImageReward [Xu et al., 2024a]. We collect the 4-step generations from each method and qualitatively compare them with the 4-step generation from the baseline VCM. As shown in Figure 6, incorporating reward feedback from any of these $\mathcal{R}_{\text{img}}$ leads to quality improvement over the baseline VCM (VC2). It is worth noting that HPSv2.1 and PickScore are fine-tuned from CLIP with human preference data. Therefore, learning from CLIP might also lead to better performance than the baseline VCM.
@@ Appendix F Qualitative Results
The prompts for the top two and bottom two rows in Figure 1 are given below:
  • With the style of low-poly game art, A majestic, white horse gallops gracefully across a moonlit beach.
  • Kung Fu Panda posing in cyberpunk, neonpunk style.